OpenStudy (anonymous):

Find the quotient: (24xy^3-16x^2y^2+32x^2y)/8xy

zepdrix (zepdrix):

\[\large \frac{24xy^3-16x^2y^2+32x^2y}{8xy}\qquad=\qquad\frac{24xy^3}{8xy}-\frac{16x^2y^2}{8xy}+\frac{32x^2y}{8xy}\] From here, we can do some division on each individual fraction.

zepdrix (zepdrix):

So for this first one, we want to: divide 24 by 8, divide x by x, and lastly, divide y^3 by y. Do you understand what happens with the y terms?

OpenStudy (anonymous):

\[\frac{ 3xy^2-2x^2y+4x^2}{x }\]

zepdrix (zepdrix):

Ok looks good so far. You just need to divide the x's now.

OpenStudy (anonymous):

soo do like we did with 8xy but with the x

zepdrix (zepdrix):

yes. It looks like you've successfully divided 8 out of each term and also divided a y out of each term. Need to do the same with the x.

OpenStudy (anonymous):

\[3y^2-2xy+4x\]

zepdrix (zepdrix):

yay good job c: